# Physics-Aware Semi-Supervised Underwater Image Enhancement

### v2

@@ Abstract
Underwater images normally suffer from degradation due to the transmission medium of water bodies. Both traditional prior-based approaches and deep learning-based methods have been used to address this problem. However, the inflexible assumption of the former often impairs their effectiveness in handling diverse underwater scenes, while the generalization of the latter to unseen images is usually weakened by insufficient data. In this study, we leverage both the physics-based underwater Image Formation Model (IFM) and deep learning techniques for Underwater Image Enhancement (UIE). To this end, we propose a novel Physics-Aware Dual-Stream Underwater Image Enhancement Network, i.e., PA-UIENet, which comprises a Transmission Estimation Steam (T-Stream) and an Ambient Light Estimation Stream (A-Stream). This network fulfills the UIE task by explicitly estimating the degradation parameters of the IFM. We also adopt an IFM-inspired semi-supervised learning framework, which exploits both the labeled and unlabeled images, to address the issue of insufficient data. Our method performs better than, or at least comparably to, eight baselines across five testing sets in the degradation estimation and UIE tasks. This should be due to the fact that it not only can model the degradation but also can learn the characteristics of diverse underwater scenes.
1 Introduction
The images captured in the underwater environment play important roles in ocean exploration. However, these images normally suffer from different degradation, due to the wavelength-dependent light absorption and light scattering caused by the underwater transmission medium. According to [1], the underwater Image Formation Model (IFM) can be expressed as:
@@ 1 Introduction
where $I^{c}(x)$ ($c\in\{R,G,B\}$) is a degraded image, $J^{c}(x)$ is the underlying clean image, $t^{c}(x)=exp(-\beta^{c}d(x))$ denotes the transmission maps, $\beta^{c}$ is the attenuation factor for each channel, $d(x)$ is the scene distance map and $A^{c}$ is the ambient light. (See Fig.1 for examples). In this context, $J^{c}(x)t^{c}(x)$ describes the distance-dependent color distortion, while $(1-t^{c}(x))A^{c}$ represents the scattering of the ambient light which decreases the contrast and lessens the visibility of the image.
@@
Figure 1: Examples of a degraded underwater image, the corresponding ambient light image $A^{c}$ and three transmission maps $t^{c}(x)$ ($c\in\{R,G,B\}$) estimated using our method, and the enhanced image produced using the estimated parameters on top of the IFM.
@@
Considering the degradation may interfere with downstream tasks, Underwater Image Enhancement (UIE) is key to underwater image processing and machine vision. Due to the complicated degradation process, this task is challenging. Many traditional methods [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] have been proposed, to address the challenge. These methods can be treated as the prior-based methods because they typically rely on a certain form of prior information. Some prior-based methods [2, 3, 4] aimed to estimate the parameters of the IFM. However, they normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene violates the prior assumption [4]. On the other hand, the other prior-based methods [9, 10, 11] directly enhance the quality of underwater images without explicitly modeling the degradation process. Due to the lack of the knowledge of underwater scenes, they cannot effectively enhance the images captured in the complex scenes and often introduce artifacts.
@@
In contrast, deep learning-based UIE approaches [14, 15, 16, 17] aim to directly learn from different types of degraded images. Therefore, the problem with the limited static prior can be alleviated. Ideally, a large number of real-world labeled training images should be utilized. However, it is difficult to collect such a large labeled underwater data set in practice. Alternatively, the best enhanced image can be selected as the reference (ground-truth) by observers from the images processed by a set of UIE algorithms [16, 18]. Nevertheless, this sort of data sets cannot represent diverse underwater scenes due to the limited size. In this case, the deep UIE model trained may overfit the training data.
@@
The synthetic images generated using the IFM [19, 14] or a Generative Adversarial Network (GAN) [20, 21, 22] were employed in order to address the problem with the lack of labeled data. However, the utilization of these images to simulate real-world underwater degradation processes is struggling. Transfer learning techniques [21, 15] were also applied for that purpose, which usually used generative models to map degraded images to the clearer images. But they are difficult to train and may produce unstable results. Recently, promising results were derived using zero-shot learning approaches [23, 24] by directly learning degradation parameters from unlabeled degraded images.
@@
To effectively model the degradation process, we propose a Physics-Aware Dual-Stream Underwater Image Enhancement Network, i.e., PA-UIENet, by injecting the physical principle into the deep neural network. This network performs the UIE task by explicitly estimating the degradation parameters of the IFM using a Transmission Estimation Steam (T-Stream) and an Ambient Light Estimation Stream (A-Stream). We further adopt an IFM-inspired semi-supervised learning framework, including a bi-directional supervised scheme and an unsupervised scheme, to overcome the limitation of insufficient data. Compared with the unsupervised methods [23, 24], the framework can be more easily trained due to the facilitation from the supervised scheme. Since the unsupervised scheme uses additional unlabeled images while the bi-directional supervised scheme exploits labeled images more effectively, it can be better trained than the supervised approaches [14, 15, 16, 17]. As a result, the generalization of the model trained using our framework is stronger than that trained using a supervised or unsupervised method. To our knowledge, this study makes the first attempt to jointly apply the physics-aware deep network and the IFM-inspired semi-supervised learning technique to the UIE task.
@@
Our main contributions can be summarized as threefold. (1) We propose a physics-aware dual-stream network for the UIE task by explicitly estimating the degradation parameters of the IFM. (2) An IFM-inspired semi-supervised learning framework is adopted, to overcome the challenge of insufficient data. (3) We conduct a series of comparative experiments on five real-world underwater testing sets along with eight baselines. These results provide the community with a set of benchmarks.
2 Related Work
2.1 Prior-Based Methods
@@
To recover clear underwater images, many methods [2, 3, 4, 5, 6, 7, 8] used the prior information to estimate the degradation parameters of the IFM. Although the Dark Channel Prior (DCP) [2] was initially proposed for image dehazing, some researchers [5, 6] employed it for underwater image enhancement, due to the similarity between the degradation processes resulted from the foggy weather and underwater environment. However, the original DCP usually yielded erroneous estimations. Therefore, more studies were performed to improve it, such as Underwater Dark Channel Prior (UDCP) [3] and Generalization of the Dark Channel Prior (GDCP) [4]. Moreover, the Haze Line Prior (HL) [7] was utilized in some studies [8, 25]. Despite these methods were designed on top of the IFM, they normally made a rigid assumption about the underwater environment, which restricted the application of these to a specific underwater scenario.
@@
On the other hand, some approaches used the more general priors [26, 9, 10, 11, 12] to enhance the quality of underwater images without taking the IFM into account, including Histogram Equalization (HE) [26, 9], Retinex-based methods [11, 12] and image fusion techniques [13, 27]. These approaches usually enhanced the contrast and produced more color-balanced results. However, they probably struggle with processing globally inhomogeneous degraded images and may introduce artifacts, such as halos and color casts, due to the lack of the knowledge of underwater scenes.
2.2 Learning-Based Methods
@@
Thanks to the powerful representation learning ability and large-scale training data [28], deep learning based methods greatly boosted the development of computer vision in both high-level [29, 30, 31] and low-level vision tasks [32, 33, 34]. However, the application of deep learning to underwater image enhancement is much less than other tasks. This dilemma should be attributed to the difficulty on collecting a large number of underwater images and the corresponding clean counterparts. Some deep learning methods [19, 35] used the fake underwater images, synthesized based on terrestrial images, to train a network for degradation parameter estimation. Since the synthesis algorithms were relatively simple, they usually could not simulate diverse underwater scenes. Besides, the content of the terrestrial images was significantly different from underwater scenes. As a result, those methods usually encountered the domain-shift problem when they were applied to real underwater images.
@@
To overcome this problem, some labeled real-world underwater image data sets, e.g., UIEB [16] and SUIM-E [18], were collected. Regarding a degraded image, a set of enhancement methods were applied. The most visually pleasant result was manually picked out as the reference of the degraded image. A UIE model can be trained using these data sets without considering the IFM. For example, the WaterNet [16] used CNNs to learn fusion weights for the purpose of fusing the enhanced results produced by the White Balance [13], HE and Gamma Correction techniques. To perform the UIE task, the UColor [17] method combined the RGB, HSV and Lab color spaces with the guidance of the IFM parameter estimated using the GDCP [4]. Since the training data sets were relatively small, these methods usually overfitted the training data and were poorly generalized to unseen images. In addition, they could not explicitly exploit the information provided by the IFM.
@@
The above-mentioned studies either are not robust to diverse underwater scenes, or lack the knowledge of underwater images. We are hence inspired to exploit both the IFM-inspired semi-supervised learning technique and the physics-aware deep neural network to address these issues.
3 Methodology
@@
Considering the importance of physical principles to the UIE task, we propose a Physics-Aware Dual-Stream Underwater Image Enhancement Network, namely, PA-UIENet. This network contains a Transmission Estimation Steam (T-Stream) and an Ambient Light Estimation Stream (A-Stream), which are used to explicitly estimate the parameters of the IFM. To address the challenge of the lack of labeled real-world underwater images, we further adopt an IFM-inspired semi-supervised learning framework, which consists of a bi-directional supervised scheme and an unsupervised scheme. Compared to the supervised or unsupervised method, the PA-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model trained is thus stronger, due to the complementary action of the two schemes.
Figure 2: The architecture of the proposed PA-UIENet, which comprises two individual streams: A-Stream and T-Stream, to estimate the degradation parameters of the IFM.
3.1 PA-UIENet
@@
As illustrated in Fig.2, the PA-UIENet contains an encoder-decoder style T-stream and a Transformer-based A-stream, used to estimate transmission maps and the ambient light respectively. Since the red channel usually encounters the more severe attenuation than the blue and green channels [1, 38], the utilization of this channel is key to color restoration. Therefore, a simple Red Channel Tuner (RCT) module is designed to adaptively emphasize the red channel before an image is fed into the T-stream. The output of the first encoder block $Enc\_1$ in the T-stream is processed by Patchify and the result is sent to the A-stream. A Residual Communication Module (RCM) connects the two blocks of the encoder of the T-stream and the A-stream at the same level, to fulfill the bi-directional feature exchange. Both the transmission maps and the ambient light, estimated using the T-stream and A-stream respectively, are fed into the IFM. As a result, the enhanced underwater image is produced.
@@
T-Stream. According to Eq. (1), the transmission map varies at different pixel locations. In this case, the network should be able to generate fine-grained representations. Hence, we build the T-stream using a CNN-based encoder-decoder network. The encoder comprises five blocks, denoted as $Enc\_1$ to $Enc\_5$, and the Max Pooling operation is applied to each block. As a result, multi-scale feature maps are derived using the encoder. Given $Enc\_5$ is used as the bottleneck, the decoder contains four blocks, denoted as $Dec\_1$ to $Dec\_4$, symmetrically. Bilinear interpolation is used to upsample the feature maps within each block. Both the encoder and decoder blocks share the same structure, which consists of two $Conv$-$BN$-$ReLU$ operations. Skip connection is used to pass the feature maps produced by an encoder block to the corresponding decoder block at the same level, which is useful for restoring the fine-grained spatial structure. The output of the last decoder block is fed into a convolutional layer. A set of transmission maps are produced in terms of different color channels.
@@
A-Stream. The ambient light describes the luminance of the scene. Since this parameter does not encode the information of a specific object, it can be treated as homogeneous. We are motivated to design the A-stream using a set of Transformer [40] blocks for the sake of leveraging the self-attention mechanism to extract the global information. Five blocks, denoted as $Trans\_1$ to $Trans\_5$, are comprised of this stream. Each block contains a self-attention layer and a two-layer Multi-layer Perceptron (MLP) unit, in which two residual connections are used respectively. Since the input of the Transformer is a sequence of visual tokens , the feature maps of $Enc\_1$ are processed using Patchify which includes the Flatten and Linear Projection operations. Considering the estimation of the ambient light does not require local characteristics and the quadratic complexity of self-attention, we downsample these maps by a factor of 16 to reduce the computational demand. The result is a set of tokens. A special token, namely, the Ambient token, is also adopted, to specifically learn the ambient light parameter. All the tokens are fed into $Trans\_1$. The Ambient token produced by the last block $Trans\_5$ is sent to a linear layer. Finally, the estimation of the ambient light is achieved.
@@
Residual Communication Module (RCM). According to the IFM, the degradation can be considered as the outcome caused by both the transmission and the ambient light. In this situation, the T-stream and A-stream will learn some common features. Therefore, the feature exchange between both the streams is likely to boost the training of each stream. Recently, the complementary action of CNNs and Transformers has been shown [41]. We are inspired to propose the RCM for the sake of bridging the two streams. Specifically, the tokens produced by a block in the A-stream are folded into a set of feature maps while the feature maps generated by a block of the encoder of the T-stream are downsampled to the shape of these maps. Then both sets of maps are concatenated along the channel dimension and are fed into a convolutional layer. Finally, the resultant maps are split into two sets along the channel dimension. Each set is added with the original feature maps of the related stream.
3.2 IFM-Inspired Semi-Supervised Learning Framework

### v3

# Semi-supervised Underwater Image Enhancement Using A Physics-Aware Triple-Stream Network

@@ Abstract
Underwater images normally suffer from degradation due to the transmission medium of water bodies. Both traditional prior-based approaches and deep learning-based methods have been used to address this problem. However, the inflexible assumption of the former often impairs their effectiveness in handling diverse underwater scenes, while the generalization of the latter to unseen images is usually weakened by insufficient data. In this study, we leverage both the physics-based Image Formation Model (IFM) and deep learning techniques for Underwater Image Enhancement (UIE). To this end, we propose a novel Physics-Aware Triple-Stream Underwater Image Enhancement Network, i.e., PATS-UIENet, which comprises a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Light Estimation Stream (A-Stream). This network fulfills the UIE task by explicitly estimating the degradation parameters of a revised IFM. We also adopt an IFM-inspired semi-supervised learning framework, which exploits both the labeled and unlabeled images, to address the issue of insufficient data. To our knowledge, such a physics-aware deep network and the IFM-inspired semi-supervised learning framework have not been used for the UIE task before. Our method performs better than, or at least comparably to, sixteen baselines across six testing sets in the degradation estimation and UIE tasks. These promising results should be due to the fact that the proposed method can not only model the degradation but also learn the characteristics of diverse underwater scenes. [^1] The models and code will be published on the acceptance of the paper.
Index Terms:
Underwater Image Enhancement (UIE), Underwater Image Processing, Image Formation Model (IFM), Deep learning, Semi-supervised Learning.
I Introduction
The images captured in the underwater environment play important roles in ocean exploration. However, these images normally suffer from different degradation, due to the wavelength-dependent light absorption and light scattering caused by the underwater transmission medium. In [1], the Image Formation Model (IFM) can be expressed as:
@@ I Introduction
where $I^{c}(x)$ ($c\in\{R,G,B\}$) is a degraded image, $J^{c}(x)$ is the underlying clean image, $t^{c}(x)=exp(-\beta^{c}d(x))$ denotes the transmission maps, $\beta^{c}$ is the attenuation factor for each channel, $d(x)$ is the scene distance map and $A^{c}$ is the ambient light. (See Fig. 1 for examples). In this context, $J^{c}(x)t^{c}(x)$ describes the distance-dependent color distortion, while $(1-t^{c}(x))A^{c}$ represents the scattering of the ambient light which decreases the contrast and lessens the visibility of the image. Despite the IFM had been widely used, its flaws in accurately modeling the degradation process was highlighted [2].
@@
Figure 1: Examples of a degraded underwater image, the corresponding ambient light image $A^{c}$ and three transmission maps $t^{c}(x)$ ($c\in\{R,G,B\}$) estimated, and the enhanced image produced using the estimated parameters of the IFM [1].
@@
Considering the degradation may interfere with downstream tasks, Underwater Image Enhancement (UIE) is key to underwater machine vision. Due to the complicated degradation process, this task is challenging. Many traditional methods [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] have been proposed, to address the challenge. These methods can be treated as the prior-based methods because they typically rely on a certain form of prior information. Some prior-based methods [3, 4, 5] aimed to estimate the parameters of the IFM. However, they normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene violates the prior assumption [5]. On the other hand, the other prior-based methods [11, 12, 13, 14, 15] directly enhance the quality of underwater images without explicitly modeling the degradation process. Due to the lack of the knowledge of underwater scenes, they cannot effectively enhance the images captured in the complex scenes and often introduce artifacts.
@@
In contrast, deep learning-based UIE approaches [18, 19, 20, 21, 22, 23, 24, 25, 26] aim to directly learn from different types of degraded images. Therefore, the problem with the limited static priors can be alleviated. Ideally, a large number of real-world labeled training images should be utilized. However, it is difficult to collect such a large labeled underwater data set in practice. Alternatively, the best enhanced image can be selected as the reference (ground-truth) by observers from the images processed by a set of UIE algorithms [22, 27]. Nevertheless, this sort of data sets cannot represent diverse underwater scenes due to the limited size. In this case, the deep UIE model trained may overfit the training data.
@@
The synthetic images generated using the IFM [28, 18] or a Generative Adversarial Network (GAN) [29, 30, 31, 20, 21] were employed in order to address the problem with the lack of labeled data. However, the utilization of these images to simulate real-world underwater degradation processes is struggling. Transfer learning techniques [30, 19] were also applied for that purpose, which usually used generative models to map degraded images to the clearer images. But they are difficult to train and may produce unstable results. Recently, promising results were derived using zero-shot learning approaches [32, 33] by directly learning degradation parameters from unlabeled degraded images.
To effectively model the degradation process, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network, i.e., PATS-UIENet, by injecting the physical principle into the deep neural network. This network performs the UIE task by explicitly estimating the degradation parameters of a revised IFM [2] using a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Estimation Stream (A-Stream). These parameters are fed into the revised IFM and the result is the image that the PATS-UIENet enhances.
@@
We further adopt an IFM-inspired semi-supervised learning framework, including a bi-directional supervised scheme and an unsupervised scheme, to overcome the limitation of insufficient data. Compared with the unsupervised methods [32, 33], the framework can be more easily trained due to the facilitation from the supervised scheme. Since the unsupervised scheme uses additional unlabeled images while the bi-directional supervised scheme exploits labeled images more effectively, it can be better trained than the supervised approaches [18, 19, 22, 23, 24, 25, 26, 34]. As a result, the generalization of the model trained using our framework is stronger than that trained using a supervised or unsupervised method.
To our knowledge, this study makes the first effort to jointly apply the physics-aware deep network and the IFM-inspired semi-supervised learning technique to the UIE task. Our main contributions can be summarized as threefold.
  • We propose a physics-aware triple-stream network for the UIE task by explicitly estimating the degradation parameters of the revised IFM. Such an IFM-motivated network has not been explored for the UIE task before.
  • An IFM-inspired semi-supervised learning framework is adopted, to overcome the challenge of insufficient data. This framework exploits the advantages of both the supervised and unsupervised methods.
@@
  • We conduct a series of comparative experiments on five real-world underwater testing sets along with sixteen baselines. These results provide the community with a set of benchmarks.
The remainder of this paper is organized as follows. The related literature is reviewed in Section II. We introduce the proposed method in Section III. The experimental settings and results are reported in Sections IV and V respectively. Finally, our conclusion is drawn in Section VI.
II Related Work
II-A Prior-Based Methods
@@
To recover clear underwater images, many methods [3, 4, 5, 6, 7, 8, 10] used the prior information to estimate the degradation parameters of the IFM. Although the Dark Channel Prior (DCP) [3] was initially proposed for image dehazing, some researchers [6, 7] employed it for underwater image enhancement, due to the similarity between the degradation processes resulted from the foggy weather and underwater environment. However, the original DCP usually yielded erroneous estimations. Therefore, more studies were performed to improve it, such as Underwater Dark Channel Prior (UDCP) [4] and Generalization of the Dark Channel Prior (GDCP) [5]. Moreover, the Haze Line Prior (HL) [8] was utilized in some studies [10, 35]. Despite these methods were designed on top of the IFM, they normally made a rigid assumption about the underwater environment, which restricted the application of them to a specific underwater scenario.
@@
On the other hand, some approaches used the more general priors [36, 11, 12, 13, 16, 14, 15] to enhance the quality of underwater images without taking the IFM into account, including Histogram Equalization (HE) [36, 11], Retinex-based methods [13, 16] and image fusion techniques [17, 37]. These approaches usually enhanced the contrast and produced more color-balanced results. However, they probably struggle with processing globally inhomogeneous degraded images and may introduce artifacts, such as halos and color casts, due to the lack of the knowledge of underwater scenes.
II-B Learning-Based Methods
@@
Thanks to the powerful representation learning ability and large-scale training data [38], deep learning-based methods greatly boosted the development of computer vision in both high-level [39, 40, 41] and low-level vision tasks [42, 43, 44]. However, the application of deep learning to underwater image enhancement is much less than other tasks. This dilemma should be attributed to the difficulty on collecting a large number of underwater images and the corresponding clean counterparts. Some deep learning methods [28, 45] used the fake underwater images, synthesized based on terrestrial images, to train a network for degradation parameter estimation. Since the synthesis algorithms were relatively simple, they usually could not simulate diverse underwater scenes. Besides, the content of the terrestrial images was significantly different from underwater scenes. As a result, those methods usually encountered the domain-shift problem when they were applied to real underwater images.
@@
To overcome this problem, some labeled real-world underwater image data sets, e.g., UIEB [22] and SUIM-E [27], were collected. Regarding a degraded image, a set of enhancement methods were applied. The most visually pleasant result was manually picked out as the reference of the degraded image. A UIE model can be trained using these data sets without considering the IFM. For example, the WaterNet [22] used CNNs to learn fusion weights for the purpose of fusing the enhanced results produced by the White Balance [17], HE and Gamma Correction techniques. To perform the UIE task, the UColor [23] method combined the RGB, HSV and Lab color spaces with the guidance of the IFM parameters estimated using the GDCP [5]. Since the training data sets were relatively small, these methods usually overfitted the training data and poorly generalized to unseen images. In addition, they could not explicitly utilize the information provided by the IFM.
@@
The above-mentioned studies either are not robust to diverse underwater scenes, or lack the knowledge of underwater images. We are hence motivated to exploit both the IFM-inspired semi-supervised learning technique and the physics-aware deep neural network, to address these issues.
III Methodology
@@
Considering the importance of physical principles to the UIE task, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network (PATS-UIENet). This network contains a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Light Estimation Stream (A-Stream), which are used to explicitly estimate the parameters of the revised IFM [2]. To address the challenge of the lack of labeled real-world underwater images, we further adopt an IFM-inspired semi-supervised learning framework, which consists of a bi-directional supervised scheme and an unsupervised scheme. Compared to the supervised or unsupervised method, the PATS-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model trained is thus stronger, due to the complementary action of the two schemes.
III-A Revised IFM
As pointed out by Akkaynak and Treibitz [2], the original IFM [1] ignored the dependencies of the backscatter coefficient on the ambient light and the optical property of water bodies. They assumed that the attenuation coefficients of the direct signal and the backscattered signal are the same. In this case, the original IFM [1] cannot describe the degradation process of underwater images adequately.
Akkaynak and Treibitz [2] further proposed a revised IFM, which can be expressed as:
$I^{c}(x)=J^{c}(x)e^{-\beta_{c}^{D}(\mathbf{D})\cdot z}+\left(1-e^{-\beta_{c}^{B}(\mathbf{B})\cdot z}\right)A^{c},$ (2)
where $\textbf{D}=\{z,\rho,E,S_{c},\beta\}$ and $\textbf{B}=\{E,S_{c},b,\beta\}$ are two sets of parameters which affect the attenuation factors of the direct signal $\beta_{c}^{D}$ and the backscatter signal $\beta_{c}^{B}$, respectively, $z$ represents the distance between the scene and the camera, $\rho$ is the reflectance spectrum of the object, $E$ is the ambient light at a certain distance, $S_{c}$ is the camera response function, $b$ is the beam scattering coefficient and $\beta$ is the beam attenuation coefficient. The above formula can be written in the more compact form as:
$I^{c}(x)=J^{c}(x)t_{D}^{c}(x,\textbf{D})+\left(1-t_{B}^{c}(x,\textbf{B})\right)A^{c}.$ (3)
It has been demonstrated that the attenuation factors of the direct signal and the backscatter signal are different while the attenuation factor of the backscatter signal is affected by the ambient light more severely [2]. For more details, please refer to [2].
III-B PATS-UIENet
@@
As illustrated in Fig. 2, the PATS-UIENet contains two encoder-decoder style streams with the same structure, i.e., D-stream and B-stream, and a Transformer-based A-stream. The three streams are used to estimate the direct signal transmission map, the backscatter signal transmission map and the ambient light, respectively. Since the red channel usually encounters more severe attenuation than the blue and green channels [1, 48], the utilization of this channel is key to color restoration. Therefore, a simple Red Channel Tuner (RCT) module is designed to adaptively emphasize the red channel before an image is fed into the transmission estimation steams.
The output of the first encoder block $Enc\_1$ in the B-stream is processed by Patchify and the result is sent to the A-stream. A Residual Communication Module (RCM) connects the two blocks of the encoder of the B-stream and the A-stream at the same level, to fulfill the bi-directional feature exchange. The two sets of transmission maps estimated using the D-stream and B-stream, respectively, and the ambient light estimated using the A-stream are fed into the revised IFM [2]. As a result, the enhanced underwater image is produced.
Figure 2: The architecture of the proposed PATS-UIENet, which comprises three individual streams, namely, D-Stream, B-Stream and A-Stream, to estimate the degradation parameters of the revised IFM [2].
@@
D-Stream. According to Eq. (3), the transmission map varies at different pixel locations. In this case, the network should be able to generate fine-grained representations. Hence, we build the D-Stream using a CNN-based encoder-decoder network. The encoder comprises five consecutive blocks, denoted as $Enc\_1$ to $Enc\_5$, and the Max-pooling operation is applied to each block. As a result, multi-scale feature maps are derived using the encoder. Given $Enc\_5$ is used as the bottleneck, the decoder contains four blocks, denoted as $Dec\_1$ to $Dec\_4$, symmetrically. Bilinear interpolation is used to upsample the feature maps within each block. Both the encoder and decoder blocks share the same structure, which consists of two $Conv$-$BN$-$ReLU$ units. Skip connection is used to pass the feature maps produced by an encoder block to the corresponding decoder block at the same level, which is useful for restoring the fine-grained spatial structure. The output of the last decoder block is fed into a convolutional layer. A set of transmission maps are produced in terms of different color channels.
B-Stream. Similarly, the B-stream is also designed as a CNN-based encoder-decoder network with the same architecture as the D-stream, motivated by Eq. (3). Unlike the D-stream, however, the output of the first encoder block $Enc\_1$ in the B-stream undergoes the Patchify processing, which includes both the flattening and linear projection operations, and the result is then passed to the A-stream. Besides, a Residual Communication Module (RCM) connects the $Enc\_2$ to $Enc\_5$ in the B-stream with the $Trans\_2$ to $Trans\_5$ in the A-stream at the same level, respectively.
@@
A-Stream. The ambient light describes the luminance of the scene. Since this parameter does not encode the information of a specific object, it can be treated as homogeneous. We are motivated to design the A-stream using a set of Transformer [50] blocks for the sake of leveraging the self-attention mechanism to extract the global information. Five blocks, denoted as $Trans\_1$ to $Trans\_5$, are comprised of this stream. Each block contains a self-attention layer and a two-layer Multi-layer Perceptron (MLP) unit, in which two residual connections are used separately.
Since the input of the Transformer is a sequence of visual tokens, the feature maps of $Enc\_1$ are processed using Patchify. Considering that the estimation of the ambient light does not require local characteristics and the quadratic complexity of self-attention is unavoidable, we downsample these maps by a factor of 16 to reduce the computational demand. The result is a set of tokens. A special token, namely, the Ambient token, is also adopted, to specifically learn the ambient light parameter. All the tokens are fed into $Trans\_1$. The Ambient token produced by the last block $Trans\_5$ is sent to a linear layer. Finally, the estimation of the ambient light is achieved.
@@
Residual Communication Module (RCM). According to the revised IFM [2], the degradation can be considered as the outcome caused by both the transmission and the ambient light. In this situation, the backscatter signal transmission estimation steam and the ambient light estimation stream will learn some common features. Therefore, the feature exchange between both the streams is likely to boost the training of each stream. Recently, the complementary action of CNNs and Transformers has been shown [51]. We are inspired to propose the RCM for the sake of bridging the two streams. Specifically, the tokens produced by a block in the A-stream are folded into a set of feature maps while the feature maps generated by a block of the encoder of the B-stream are downsampled to the shape of these maps. Then both sets of maps are concatenated along the channel dimension and are fed into a convolutional layer. Finally, the resultant maps are split into two sets along the channel dimension. Each set is added with the original feature maps of the related stream.
III-C IFM-Inspired Semi-supervised Learning Framework